OpenStudy (anonymous):

Find an equation of the line tangent to the graph of y=3/x^4 at the point (5, 3/625).

OpenStudy (anonymous):

we know the slope of the tan line at a given point is the first derivative... can you find the first derivative of this function?

OpenStudy (anonymous):

If I'm right, I believe it's -12x^-3! Maybe!

OpenStudy (anonymous):

to the -5 actually, my apologies!

OpenStudy (anonymous):

find the derivitive of f(x) call it f'(x) plug in 5 f'(5) and that is your answer.

OpenStudy (anonymous):

okay, can you confirm if the derivative I have no is right? I currently have -12x^-5

OpenStudy (anonymous):

-12/x^5 yes it is right

OpenStudy (anonymous):

so just -12/(5)^5

OpenStudy (anonymous):

but I need a y=(answer) will this still work?

OpenStudy (anonymous):

y=mx+b 3/625=(-12/3125)*5+b sfb 3/125 y=(-12/3125)x+3/125

OpenStudy (anonymous):

YES! THANK YOU SO MUCH! I OWE YOU MY LIFE!

OpenStudy (anonymous):

aww shucks :)